OpenStudy (anonymous):

Help Simplifying 7 + 4(2x - 3) + 2x

OpenStudy (anonymous):

umm first do 4(2x-3) which is 8x-12 than u add 7 and 2x which is 10x-5 and also u can simplify by dividing the whole term by 5 so its 2x-1

OpenStudy (austinl):

PEMDAS THE ORDER IN WHICH YOU EVALUATE: Parentheses Exponents Multiplication Division Addition Subtraction

OpenStudy (anonymous):

i think its 10x-5 or 2x-1

OpenStudy (anonymous):

7+4(2x-3)+2x You apply the distributive property to 4(2x-3). So you do 4 times 2x which equals to 8x. Then you do 4 times (-3) which equals (-12). So re write the equation with the simplified answers. 7+8x-12+2x Now you have 2 terms of "x's" so you can combine 8x and 2x. 7+10x-12 then you do 7+(-12) you get -5 FINAL ANSWER: 10X-5

OpenStudy (austinl):

That is correct.

OpenStudy (anonymous):

I also found this on the internet: 7 + 4(2x - 3) + 2x Multiply x and 2 Multiply x and 1 The x just gets copied along. The answer is x x 2*x evaluates to 2x 2*x-3 evaluates to 2x-3 Multiply 4 by 2x-3 we multiply 4 by each term in 2x-3 term by term. This is the distributive property of multiplication. Multiply 4 and 2x Multiply 1 and x The x just gets copied along. x 4 × 2x = 8x Multiply 4 and -3 1 4 × -3 = -12 4*(2*x-3) evaluates to 8x-12 7 + -12 = -5 The answer is -5+8x 7+4*(2*x-3) evaluates to -5+8x Multiply x and 2 Multiply x and 1 The x just gets copied along. The answer is x x 2*x evaluates to 2x 8x + 2x = 10x The answer is 10x-5 7+4*(2*x-3)+2*x evaluates to 10x-5 The final answer is 10x-5